# Towards Scalable Automated Alignment of LLMs: A Survey

### v1

@@ 5.1.5 Filtering SFT Data using Reward Model
High-quality SFT plays a crucial role in enhancing LLM performance (Zhou et al., 2023a), therefore some studies use reward models to filter the training data. The main paradigms are categorized into learning from best response and learning from ranked results. Learning from the best response is often referred to as Best of N or Reject sampling (Touvron et al., 2023; Yuan et al., 2023b). This approach typically involves using a reward model to choose the high-quality data among multiple responses for refining the model (Dong et al., 2023). In addition to learning from top responses, the LLM can also learn from ranked data. Yuan et al. (2023a) propose Rank Responses to align Human Feedback (RRHF) to align LLM using ranking loss. Lu et al. (2022) propose using a reward model to grade data based on their scores and employing various reward tokens to regulate their generation. This approach helps prevent the learning of undesirable behaviors.
@@ 5.2 Binary Verifier
Even if the final answer is correct, there may still be errors in the reasoning process, limiting the effectiveness of the outcome verifier. To address this problem, the process verifier is employed to assess the correctness of reasoning step for more detailed verification (Lightman et al., 2023; Uesato et al., 2022). The process verifier can be used to train a more effective reasoner (Ying et al., 2024; Shao et al., 2024). Inspired by the human reasoning mechanism, Zhu et al. (2023b) propose Cooperative Reasoning (CoRe) to produce synthesized training data for reasoning where process verifier is used to generate the feedback of the generation of generator. Many studies are dedicated to training the verifier using automatically generated data due to the difficulty of collecting step-wise supervised signals.  Wang et al. (2024e) and Wang et al. (2024h) train the process verifier with automatically constructed data that is collected by Monte Carlo Sampling. Moreover, the process verifier is can be applied in decoding to select the correct reasoning path (Khalifa et al., 2023). Some studies focus on how to complete the final reasoning path efficiently. Ma et al. (2023b) propose a heuristic greedy search algorithm based on the verifier’s feedback. Li et al. (2023c) propose using the verifier to filter the reasoning steps generated using diverse prompts.
Binary verifier is an important way to achieve automated alignment in objective tasks such as math. However, training the verifier, especially the process verifier is very difficult and requires a large amount of annotated data. Therefore, in the future, in order to further achieve automated alignment, researchers can focus on how to automatically construct a process verifier.
@@ 5.3 Text Critic
Text signals contain more semantics than scalar and binary signals, enabling models to intuitively align with humans. By integrating text feedback, LLMs can improve their output alignment with humans. These refined outputs can then be used as supervised data for further aligning the LLMs (Scheurer et al., 2022; Chen et al., 2024a). Text signals generated by the critique model have shown the potential to improve the outputs of LLM, with feedback typically obtained through prompting LLMs. The text critic can be other LLMs (such as GPT-4) (Koutcheme et al., 2024; An et al., 2024) or LLM itself (i.e., self-critique) (Saunders et al., 2022; Wang et al., 2023d). Since the self-critique of LLM remains a challenge (Luo et al., 2023b), text signals for alignment are still underexplored.
@@ 6 Aligning Through Environment Feedback
  • Human collective intelligence (§ 6.2), where models receive judgements from human societies to calibrate their internal values.
@@
Figure 8: Alignment through various types of environment feedback by social interactions, human collective intelligence, signals from tool execution or embodied environment.
@@ 6.1 Social Interactions
For instance, Stable Alignment (Liu et al., 2023a) draws inspiration from how humans learn to navigate social norms and reach consensus on value judgments of societal issues through discussions. They introduce an LLM-based multi-agent framework to simulate social interactions within human society, highlighted by a three-tiered method comprising imitation, self-critic and realignment. Alignment datasets are extracted during the simulation process, correspondingly. Wang et al. (2024f) extend this method to multi-turn interactions under realistic social scenarios. They propose an interactive learning method to teach language agents to reach goals through social interactions among characters playing different social roles. Building upon the same methodology of simulated social interactions, other research works explore the intricacies of moral discussions (Sun et al., 2023a), political debates (Taubenfeld et al., 2024) and task-oriented dialogue (Ulmer et al., 2024) wherein AI systems are guided by some predefined discussion schema.
6.2 Human Collective Intelligence
Another way compensate to simulated social interactions involves resorting to crowd-sourcing effort from human. However, distinct from previous crowd-sourcing annotations in traditional NLP or RLHF data collection, alignment signals are collected in more primitive forms to achieve broader scalability of supervision and oversight. These signals are often termed as norms (Ammanabrolu et al., 2022), Rules of Thumb (RoTs) (Forbes et al., 2020; Ziems et al., 2022), or Constitutions (Bai et al., 2022b) in different contexts. Through such process, consensus on what basic principles AI systems should follow could be achieved among groups of people, and collective intelligence can be reached by synthesizing various human ideas on values into a target for aligning models (Klingefjord et al., 2024).
Toward this goal, current works are carried out to mainly solve two aspects of research questions. One of the questions is how to extract alignment signals from human values. For this question, Constitutional AI (CAI) (Bai et al., 2022b) provides a basic attempt to convert the AI Constitutions into supervision and preference datapoints. In their setting, human supervision relies entirely on a set of principles governing AI behavior, called Constitutions. A small number of examples are used for few-shot prompting LLMs to create SFT data or preference feedback according to those constitutional principles.
@@ 6.2 Human Collective Intelligence
The other question is how to elicit values from human public more effectively. For example, in a program titled Democratic inputs to AI555openai.com/blog/democratic-inputs-to-ai, an open call for the design of prototype systems to promote the democratic process of AI was issued. Specifically, it addresses enabling a representative cross-section of people to exchange views and ultimately reach a consensus on the rules AI systems should follow. Subsequently, in the Collective Constitutional AI (CCAI) project (Anthropic, 2023), researchers design questionnaire to ask participants to draft a constitution for AI system. Through the way of collective constitution drafting, participants can assess the existing constitutional principles, rate their acceptance scores, and propose new ones they believe AI should adhere to.
@@ 6.3 Tool Execution Feedback
Tools’ execution feedback can serve as additional signals beyond human-labor corpus to teach models better learn tasks. Code generation tasks provide such examples. CodeRL (Le et al., 2022) conducts unit tests by code compiler to receive feedback signals and use them to train a critic model, with which the code generation model is further trained using deep reinforcement learning. Self-debugging (Chen et al., 2023d) designs an interactive code debugging pipeline by adding a code explanation stage beside the code execution feedback, where large language models are asked to debug its own generated code given these feedback information. Similarly, SelfEvolve (Jiang et al., 2023) receives error message from interpreter and refine the answer code based on this feedback.
Beyond code generation, there are also studies designing unified frameworks to align language models with the execution feedback from multiple other sources. CRITIC (Gou et al., 2024) utilizes a verify-then-correct process to obtain external feedback from various external tools, including search engines, code interpreters and text APIs, thus enabling language models correcting the outputs with tool-interactive critiquing. Taking a step further, Wang et al. (2024a) augment LLMs with a dynamic memory mechanism, which allows the LLMs to progressively learn how to accurately use tools. Qiao et al. (2024) propose Reinforcement Learning with Execution Feedback to reinforce LLMs with execution results of tools.
@@ 6.4 Embodied Environment
Embodied intelligence involves agents that perceive signals and take actions within some physical environments, requiring not only language understanding but also abilities of reasoning and decision making in a space with large number of states (Roy et al., 2021). Leveraging large language models in embodied AI settings represents a compelling endeavor which can bring benefits from two aspects. On one hand, the powerful generalization ability of large language model can be utilized to facilitate action plan in natural language without preliminary learning. On the other hand, commonsense knowledge and physical understanding ability of large language models can be aligned with feedback signals from environment.
Recent studies show the potential of LLMs to be deployed in robotics for real-world interactions while also pose challenges for grounding LLMs to understand the physical world (Wang et al., 2023a; Ahn et al., 2022). To better align language models which have not been pre-tarined on those embodied environment, Xiang et al. (2023) adopt an embodied simulator as world model to provide feedback signals when LLMs interact with objects and execute actions in this environment with or without a task goal. Then the explored world states would be collected as embodied experiences to fine-tune the LLM offline subsequently. In parallel, other studies utilize online reinforcement learning (Carta et al., 2023; Tan et al., 2024) to learn or iterative offline learning framework (Song et al., 2024b) to update policy for embodied agent through trial and error in the interactive environments.
@@ 6.5 Discussion
In this section, we reviewed some current studies that can be uniformly observed from the perspective of aligning through environment feedback. While most of the works target at specific downstream applications, actively learning from environment has always been a key pursue in artificial intelligence. Despite their valuable contributions, current methods still show limitations and leave open problems for future research. The main limitation arises from the gap between simulated environments and the real world.
Most of the current works derive alignment signals from simulated environments, while feedback from the real environment can be noisy or ambiguous. For example, in aligning through social interaction, Liu et al. (2023a) point out that current studies usually assume a static view of social norms, overlooking their dynamic and evolving natures. In fact, social values are complex, context-dependent, and sometimes contradictory. Encoding these values into AI systems in a comprehensive and coherent manner is a significant challenge. Thus, how to effectively bridge these changeable and divergent social values is of critical importance in alignment with human society.
In addition, simulated environments can be oversimplified and cannot fully catch the complexity of real-world, which challenges the robustness and generalization of aligned models. As for the embodied environment track of works, research are often conducted within sandbox environments and different settings are mutually independent. The action spaces of these simulated environments are still limited (Tan et al., 2024) compared to the real environment with potentially infinite degree of freedom. As a result, models trained in one setting might not generalize well to others, and ensuring models maintain alignment across different environments remains a problem.
In conclusion, the topic of integrating environmental feedback into AI alignment open up a promising and emerging track, while significant research problems remain to be explored. Future research can focus on bridging the gap between simulated and real environments to develop more adaptive, reliable, and ethically aligned AI.
@@ 7 Underlying Mechanism of Automated Alignment
  • RQ2: Why does Self-feedback work? Self-feedback is widely applied across various paradigms of automated alignment, such as serving as an inductive bias in section 3.1.2, constructing data for preference learning in section 5 and acting as a crucial technique for automated iterative alignment (Yuan et al., 2024; etc.). Understanding why does it work is vital for comprehending and enhancing all paradigms involving it.
@@ 7.2 Why does self-feedback work?
Feedback capability refers to the capacity to provide information or guidance to a given input based on specific standards. As we discussed above, such capability is extensively utilized in various paradigms of automated alignment. For example, in RLAIF, LLM itself substitutes for human assessment to construct preference data (Yuan et al., 2024; etc.). Additionally, LLM can continuously optimize its output based on self-feedback, a process known as self-refine (Bai et al., 2022b; Madaan et al., 2023; Tan et al., 2023; etc.). Nonetheless, there is much debate about whether and why LLMs can provide effective feedback to their own responses. In the following, we systematically summarize each representative viewpoint.
@@
Li et al. (2024f) and Lin et al. (2024b) suggest that models possess certain knowledge that cannot be directly utilized for generation but can be employed for providing feedback. Li et al. (2024f) also discover significant Generator-Validator inconsistency in LLMs when generating and validating answers. Similarly, Lin et al. (2024b) find that LLMs possess substantial knowledge that cannot be expressed through generation and correction but can be utilized for critique. Furthermore, other studies (Liu et al., 2024a; Li et al., 2024e) argue that feedback capability is a byproduct of the model’s ability to follow instructions. Therefore, during the alignment process, the model’s feedback capacity improves alongside its ability to follow instructions, a phenomenon supported by experimental evidence (Liu et al., 2024a; Li et al., 2024e). However, some other studies posit that the feedback capability of LLM is illusory, suggesting that it may rely on specific data and exhibit biases. West et al. (2023) note that while current LLMs gain generative capabilities to replicate expert-level outputs through training, they lack understanding capacities relevant to critique. Moreover, Huang et al. (2023b) observe that for reasoning tasks, LLMs even experience performance degradation after self-correction at times because that they cannot properly judge the correctness of their reasoning. Additionally, Zheng et al. (2024b) find that although powerful LLMs like GPT-4 achieve high agreement with human evaluators, similar to human-human agreement levels, LLM-based evaluation face various challenges such as position bias (Zheng et al., 2024b; Wang et al., 2023c), verbosity bias (Zheng et al., 2024b; Wu and Aji, 2023), self-enhancement bias (Zheng et al., 2024b; Liu et al., 2023b), as well as limited capability in certain scenarios such as grading math, reasoning questions (Zheng et al., 2024b), high-quality summaries (Shen et al., 2023a) and so on (Lan et al., 2024). Researchers propose many methods to further improve the capability of self-feedback, such as meta critic (Sun et al., 2024a), autocalibrate (Liu et al., 2023d), split and merge (Li et al., 2023e) and using external tools (such as search engines, code interpreters, etc.) to cross-check (Gou et al., 2024), thus refining their initial response, etc.

### v2

@@ 5.1.5 Filtering SFT Data using Reward Model
High-quality SFT plays a crucial role in enhancing LLM performance (Zhou et al., 2023a). Therefore some studies use reward models to filter the training data. The main paradigms are categorized into learning from best response and learning from ranked results. Learning from the best response is often referred to as Best of N or Reject sampling (Touvron et al., 2023; Yuan et al., 2023b). This approach typically involves using a reward model to choose the high-quality data among multiple responses for refining the model (Dong et al., 2023). In addition to learning from top responses, the LLM can also learn from ranked data. Yuan et al. (2023a) propose Rank Responses to align Human Feedback (RRHF) to align LLM using ranking loss. Lu et al. (2022) propose using a reward model to grade data based on their scores and employing various reward tokens to regulate their generation. This approach helps prevent the learning of undesirable behaviors.
@@ 5.2 Binary Verifier
Even if the final answer is correct, there may still be errors in the reasoning process, limiting the effectiveness of the outcome verifier. To address this problem, the process verifier is employed to assess the correctness of reasoning step for more detailed verification (Lightman et al., 2023; Uesato et al., 2022). The process verifier can be used to train a more effective reasoner (Ying et al., 2024; Shao et al., 2024). Inspired by the human reasoning mechanism, Zhu et al. (2023b) propose Cooperative Reasoning (CoRe) to produce synthesized training data for reasoning where process verifier is used to generate the feedback of the generation of generator. Many studies are dedicated to training the verifier using automatically generated data due to the difficulty of collecting step-wise supervised signals.  Wang et al. (2024e) and Wang et al. (2024h) train the process verifier with automatically constructed data that is collected by Monte Carlo Sampling. Moreover, the process verifier can be applied in decoding to select the correct reasoning path (Khalifa et al., 2023). Some studies focus on how to complete the final reasoning path efficiently. Ma et al. (2023b) propose a heuristic greedy search algorithm based on the verifier’s feedback. Li et al. (2023c) propose using the verifier to filter the reasoning steps generated using diverse prompts.
Binary verifier is important for achieving automated alignment in objective tasks such as math. However, training the verifier, especially the process verifier is very difficult and requires a large amount of annotated data. Therefore, in the future, in order to further achieve automated alignment, researchers can focus on how to automatically construct a process verifier.
@@ 5.3 Text Critic
Text signals contain more semantics than scalar and binary signals, enabling models to intuitively align with humans. By integrating text feedback, LLMs can improve their output alignment with humans. These refined outputs can then be used as supervised data for further aligning the LLMs (Scheurer et al., 2022; Chen et al., 2024a). Text signals generated by the critique model have shown the potential to improve the outputs of LLMs, with feedback typically obtained through prompting LLMs. The text critic can be other LLMs (such as GPT-4) (Koutcheme et al., 2024; An et al., 2024) or LLM itself (i.e., self-critique) (Saunders et al., 2022; Wang et al., 2023d). Since the self-critique of LLM remains a challenge (Luo et al., 2023b), text signals for alignment are still underexplored.
@@ 6 Aligning Through Environment Feedback
  • Human Shared-Values (§ 6.2), where models receive judgements from human societies to calibrate their internal values.
@@
Figure 8: Alignment through various types of environment feedback by social interactions, human shared-values, signals from tool execution or embodied environment.
@@ 6.1 Social Interactions
For instance, Stable Alignment (Liu et al., 2023a) draws inspiration from how humans learn to navigate social norms and reach consensus on value judgments of societal issues through discussions. They introduce an LLM-based multi-agent framework to simulate social interactions within human society, highlighted by a three-tiered method comprising imitation, self-critic and realignment. Alignment datasets are extracted during the simulation process, correspondingly. Wang et al. (2024f) extend this method to multi-turn interactions under realistic social scenarios. They propose an interactive learning method to teach language agents to reach goals through social interactions among characters playing different social roles. Pang et al. (2024) introduce social scene simulation which employs LLM to play different social roles in scenario relative to the instruction, enabling the model to take social consequences behind the instruction into account to revise its initial response accordingly. Building upon similar methodology of simulated social interactions, other works explore the intricacies of moral discussions (Sun et al., 2023a), political debates (Taubenfeld et al., 2024) and task-oriented dialogue (Ulmer et al., 2024) wherein AI systems are guided by some predefined discussion schema.
6.2 Human Shared-Values
Another way compensate to simulated social interactions involves resorting to collective efforts from human to derive principles that AI should align to. Currently, in this line of works, alignment signals are usually termed as norms (Ammanabrolu et al., 2022), Rules of Thumb (RoTs) (Forbes et al., 2020; Ziems et al., 2022), or Constitutions (Bai et al., 2022b) in different contexts, which provide greater scalability compared to the data-point-level annotations in traditional NLP and RLHF data collections. In Constitutional AI (CAI) (Bai et al., 2022b), researchers have designed a set of AI Constitutions. However, such in-house constitutions compiled by members from a research team fall short to broadly represent values of different groups of people.
@@ 6.2 Human Shared-Values
In order to achieve consensus on what basic principles AI systems should follow among groups of people and reach human shared-values by synthesizing various human ideas into a set of target for model alignment, follow-up efforts are dedicated to enable broader public to collectively shape the behavior of AI systems. For example, in a program titled Democratic inputs to AI444openai.com/blog/democratic-inputs-to-ai, an open call for the design of prototype systems to promote the democratic process of AI was issued. Specifically, it addresses enabling a representative cross-section of people to exchange views and ultimately reach a consensus on the rules AI systems should follow. Subsequently, in the Collective Constitutional AI (CCAI) project (Anthropic, 2023; Huang et al., 2024), researchers establish a multi-stage process for sourcing and integrating public input into LMs based on questionnaire. Through the way of collective constitution drafting, participants can assess the existing constitutional principles, rate their acceptance scores, and propose new constitutions they believe AI should be better to adhere to, thus reduce the bias of aligned LLMs. Beyond these pioneering research, large language models themselves are also involved to help conclude shared-values from human crowds more efficiently. Klingefjord et al. (2024) propose Moral Graph Elicitation (MGE) to elicit and reconcile values from human participants, which utilizes an LLM to interview participants about their values in particular contexts.
@@ 6.3 Tool Execution Feedback
Tools’ execution feedback can serve as additional signals beyond human-labor corpus to better ground models to tool usage and reduce their hallucinations during interacting with tools. Code generation tasks provide such examples. CodeRL (Le et al., 2022) conducts unit tests by code compiler to receive feedback signals and use them to train a critic model, with which the code generation model is further trained using deep reinforcement learning. Self-debugging (Chen et al., 2023d) designs an interactive code debugging pipeline by adding a code explanation stage beside the code execution feedback, where LLMs are asked to debug its own generated code given these feedback information. Similarly, SelfEvolve (Jiang et al., 2023) receives error message from interpreter and refine the answer code based on this feedback.
In addition to code generation, there are also studies designing unified frameworks to align language models with the execution feedback from multiple other sources. CRITIC (Gou et al., 2024) utilizes a verify-then-correct process to obtain external feedback from various external tools, including search engines, code interpreters and text APIs, thus enabling language models correcting the outputs with tool-interactive critiquing. Taking a step further, Wang et al. (2024a) augment LLMs with a dynamic memory mechanism, which allows the LLMs to progressively learn how to accurately use tools. Qiao et al. (2024) propose Reinforcement Learning with Execution Feedback to reinforce LLMs with execution results of tools.
@@ 6.4 Embodied Environment
Embodied intelligence involves agents that perceive signals and take actions within some physical environments, requiring not only language understanding but also abilities of reasoning and decision making in a space with large number of states (Roy et al., 2021). Leveraging large language models in embodied AI settings represents a compelling endeavor which can bring benefits from two aspects. On one hand, the powerful generalization ability of LLM can be utilized to facilitate action plan in natural language without preliminary learning. On the other hand, commonsense knowledge and physical understanding ability of LLMs can be aligned with feedback signals from environment.
Recent studies show the potential of LLMs to be deployed in robotics for real-world interactions while also pose challenges for grounding LLMs to understand the physical world (Wang et al., 2023a; Ahn et al., 2022). To better align language models which have not been pre-tarined on those embodied environment, Xiang et al. (2023) adopt an embodied simulator as world model to provide feedback signals when LLMs interact with objects and execute actions in this environment with or without a task goal. Then the explored world states would be collected as embodied experiences to fine-tune the LLM offline subsequently. In parallel, other studies utilize online reinforcement learning (Carta et al., 2023; Tan et al., 2024) to learn or iterative offline learning framework (Song et al., 2024b) to update policy for embodied agent through trial and error in the interactive environments. Beyond aligning specialist agents in isolated environments, Xi et al. (2024) introduce a novel framework called AGENTGYM to develop generally-capable LLM-based agents by facilitating their evolution through diverse environments and tasks.
@@ 6.5 Discussion
In this section, we reviewed some current studies that can be uniformly observed from the perspective of aligning through environment feedback. While most of the works target at specific downstream applications, actively learning from environment has always been a key pursue in artificial intelligence. Despite their valuable contributions, current methods still show limitations and leave open problems for future research. One of the primary limitations arises from the gap between environment signals inspected in research and those in the real world.
On the one hand, simulated environments may be lake of representation ability to fully catch the complexity of real-world. For example, to derive social interaction signals for alignment, most of the current works are based on simulated environments, while feedback from the real environment can be more noisy or ambiguous. For example, Liu et al. (2023a) point out that current studies usually assume a static view of social norms, overlooking their dynamic and evolving natures. Similarly, for the embodied environment track of works, research are often conducted within sandbox environments and different settings are mutually independent. The action spaces of these simulated environments are still limited (Tan et al., 2024) compared to the real environment with potentially infinite degree of freedom. As a result, models trained in one setting might not generalize well to others, and ensuring models maintain alignment across different environments remains a challenge.
On the other hand, signals from human shared-values could still suffer from bias even they are linked to the real-world environment. For example, the values sampled from a small society groups can quite differ from the universally acknowledged ones. Additionally, voices on the internet may not always represent the actual values of most people in the real world and peoples’ perceptions are tend to be influenced by social media. Since alignment with human shared-values is still a nascent territory, most of current works are tentative and invite limited number of participants not more than thousands. The potential challenges of current methods when scaling to hundreds of thousands of people may be unanticipated. As a result, how to inject human values into AI systems in a comprehensive and coherent manner to achieve as much unbiased alignment as possible is still an open problem.
In conclusion, the topic of integrating environmental feedback into AI alignment opens up a promising and emerging track, while significant research problems remain to be explored. Future research can focus on bridging the gap between simulated and real environments to develop more adaptive, reliable, and unbiased AI.
@@ 7 Underlying Mechanism of Automated Alignment
  • RQ2: Why does Self-feedback work? Self-feedback is widely applied across various paradigms of automated alignment, such as serving as an inductive bias in section 3.1.2, constructing data for preference learning in section 5 and acting as a crucial technique for automated iterative alignment (Yuan et al., 2024b; etc.). Understanding why does it work is vital for comprehending and enhancing all paradigms involving it.
@@ 7.2 Why does self-feedback work?
Feedback capability refers to the capacity to provide information or guidance to a given input based on specific standards. As we discussed above, such capability is extensively utilized in various paradigms of automated alignment. For example, in RLAIF, LLM itself substitutes for human assessment to construct preference data (Yuan et al., 2024b; etc.). Additionally, LLM can continuously optimize its output based on self-feedback, a process known as self-refine (Bai et al., 2022b; Madaan et al., 2023; Tan et al., 2023; etc.). Nonetheless, there is much debate about whether and why LLMs can provide effective feedback to their own responses. In the following, we systematically summarize each representative viewpoint.
@@
Li et al. (2024g) and Lin et al. (2024b) suggest that models possess certain knowledge that cannot be directly utilized for generation but can be employed for providing feedback. Li et al. (2024g) also discover significant Generator-Validator inconsistency in LLMs when generating and validating answers. Similarly, Lin et al. (2024b) find that LLMs possess substantial knowledge that cannot be expressed through generation and correction but can be utilized for critique. Furthermore, other studies (Yuan et al., 2024b; Li et al., 2024f) argue that feedback capability is a byproduct of the model’s ability to follow instructions. Therefore, during the alignment process, the model’s feedback capacity improves alongside its ability to follow instructions, a phenomenon supported by experimental evidence (Yuan et al., 2024b; Li et al., 2024f). However, some other studies posit that the feedback capability of LLM is illusory, suggesting that it may rely on specific data and exhibit biases. West et al. (2023) note that while current LLMs gain generative capabilities to replicate expert-level outputs through training, they lack understanding capacities relevant to critique. Moreover, Huang et al. (2023b) observe that for reasoning tasks, LLMs even experience performance degradation after self-correction at times because that they cannot properly judge the correctness of their reasoning. Additionally, Zheng et al. (2024b) find that although powerful LLMs like GPT-4 achieve high agreement with human evaluators, similar to human-human agreement levels, LLM-based evaluation face various challenges such as position bias (Zheng et al., 2024b; Wang et al., 2023c), verbosity bias (Zheng et al., 2024b; Wu and Aji, 2023), self-enhancement bias (Zheng et al., 2024b; Liu et al., 2023b), as well as limited capability in certain scenarios such as grading math, reasoning questions (Zheng et al., 2024b), high-quality summaries (Shen et al., 2023a) and so on (Lan et al., 2024). Researchers propose many methods to further improve the capability of self-feedback, such as meta critic (Sun et al., 2024a), autocalibrate (Liu et al., 2023d), split and merge (Li et al., 2023e) and using external tools (such as search engines, code interpreters, etc.) to cross-check (Gou et al., 2024), thus refining their initial response, etc.